OpenStudy (anonymous):

Medal! We can use the expression ps + pt + p to calculate the total of our bill at the restaurant. p = price s = rate of sales tax t = rate of tip Your bill before tax or tip was $32 The sales tax is 7% You left a 15% tip. Which answer makes the most sense as a total with tax and tip? A. $39.04 B. $92.40 C. $82.24 D. $69.20

OpenStudy (bela):

The answer is 39.38 so I'd go with A???

OpenStudy (anonymous):

....

OpenStudy (welshfella):

just plug in the given values 15% = 0.15 and 7% = 0.07 p =32 , pt = 32*0.15 and ps = 32*0.07

OpenStudy (welshfella):

add the above to get your answer

OpenStudy (welshfella):

answer = 32 + (32*0.15) + (32*0.07)

OpenStudy (welshfella):

I wish people would not just give a direct answer...